OpenStudy (anonymous):

A spring has a spring stiffness constant, k, of 440 N/m. How much must the spring be stretched to store 25 J of potential energy?

OpenStudy (anonymous):

\[(1/2)kx ^{2}= 25 J\] \[x ^{2}= 50/k\] \[x = \sqrt{5/44}\]